Operations Report – December 4th

SOL: SOL 3

Name of person filing report: HuntJ

Non-nominal systems:

  1. Wi-Fi Modem (lack of range).  The wifi in a roughly 5 feet circle from the router, any farther and signal drops off line.  We have tested this and created a signal service map by carrying laptops around the habitat to determine when it picks up signal.  No known fix, will operate internet within this range for now.
  2. HAB Eastern Wall has sustained wind damage, panels are beginning to peel off wall. When wind blows the panels swing and hit support beam causing HAB to vibrate.  In order to fix we need HAB material adhesive delivered, after which we will go EVA and patch the damage,
  3. ATV Yamaha 250 #3 Back Right Wheel is loose, observed to be wobbling during driving.  We will assess a fix during tomorrows EVA – it is possible it simply needs to be tightened, if there is actual mechanical damage then we will move the ATV and declare it off limits.
  4. 3 broken old radios.  All newly delivered radios are functional (thank you supply drop).  Assessing broken radios to determine a fix, if none identified will mark as trash and place in a ‘broken equipment bin’.  4 EVA earpieces were declared nonfunctional, though this may have been due to incorrect settings, will continue to investigate during tomorrows EVA

Generator (hours run): 15.4 Hours

Solar— SOC % : @ 7:34 AM: 65%

                            @ 8:00 AM: 100 % (Generator stopped at this time)

@ 5:34 AM: 42% (Generator started at this time)

Diesel –  48% Full.

Propane –  79 Percent Volume .

Ethanol Free Gasoline (5 Gallon containers for ATV) –   gallons. 2.5 Gallons. One 5 gallon tank empty, second tank half full. Request for extra fuel.

Water (trailer) –  600 Gallons

Water (static) –   500 gallons (We have water again!!)

Trailer to Static Pump used –  Yes

Water (loft) – Static to Loft Pump used –  Yes

Water Meter:  35 gallons

Toilet tank emptied: No

ATV’s Used: 3 Yamaha 250’s and 1 Honda ATV used.

Oil Added? No

ATV Fuel Used:  7.5 gallons used to fill all ATVs to full, though the ATVs did not start with full tanks for our mission.

# Hours the ATVs were Used today: 02:30 hours

Notes on ATVs: See above for wobbly ATV wheel issue.

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Trailer to Static water tank was used today. Pump was dirty and needed cleaning and we did so. Actual process was nominal. We are now full on water again. When inspecting trailer tank water the front tank had a layer of FOD and detritus at the water line. We will keep watch of filter usage in case this increases our burn through rate on filters.

Summary of internet:  Wifi was fixed this morning, though range is limited (see above). Please send hardware manual for the router and modem to allow us to diagnose issues and repair anything that comes up in the future (for example – to allow us to diagnose what certain light combinations signify)

Summary of suits and radios: See above

Summary of Hab operations: Brick Communication Experiment / CPR experiments successful, exercise on all crew. EVA Outer Door repaired.

Summary of GreenHab operations: No fire extinguisher observed – please deliver in future supply drop.

Summary of ScienceDome operations: Soil Salinity experiments (previous crew), dehydrating dirt for use as counterbalance and weights in exercise equipment (filling bags and medicine balls)

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support:

Need extra fuel for ATVs. Of the two red fuel cans, one is empty, one is half full.

Pre-Mission food inventory is complete, will be sent as it’s own report. Please reference that report for requested supplies to address gaps in current inventory. Please give ETA for supply drop so we can adequately ration food.

Thank you for your assistance!

Operations Report – December 3rd

Crew 184 Operations Report  03/12/2017

SOL: SOL 2

Name of person filing report: HuntJ

Non-nominal systems: Toilet Tank, Water (static)

Notes on non-nominal systems:  Toilet Tank: Tank was particularly filled up during purge and caused all crew to have to move upstairs for duration of purge. Workaround is to purge every other day of ops (e.g Sol 4 is next purge) to make upkeep easier on all.

We also request a resupply of the TB-Cide Quat Cleaner Deordorizer/Disinfectant. We are out after the last purge.

 

                                                    Water (Static): Water is at 50 Gallons which is alarmingly low. Our trailer water is at 1000 gallons but we need the means to transfer the water. Request to get Trailer water moved to Static water or permission and high level instructions to do it ourselves.

Generator (hours run): 13.2 Hours

Solar— SOC % (Before generator is run at night) Not nightfall as of submission, but will report this info in the next report.

Diesel –  50% full

Propane –  Did not check (No EVA), assume to be 79 percent volume due to past readings and trends.

Ethanol Free Gasoline (5 Gallon containers for ATV) –   gallons. 5 gallons

Water (trailer) –   1000 Gallons

Water (static) –   50 Gallons

Trailer to Static Pump used –  No

Water (loft) – Static to Loft Pump used –  Yes

Water Meter:  20 gallons (46 once Pump used)

Toilet tank emptied: Yes

ATV’s Used: None

Oil Added? No

ATV Fuel Used:  None

# Hours the ATVs were Used today: 00

Notes on ATVs: N/A

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments:

Summary of internet:  Still good so far.

Summary of suits and radios: All nominal

Summary of Hab operations: Static Water Tank almost empty, need means to transfer Trailer water over to Static.

Summary of GreenHab operations:  All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Is there a way we can check to see how much internet we are using during the day? We tried the link provided in a different log template but it didn’t take us to anything.

Operations Report – December 2nd

Crew 184 Operations Report 02/12/2017

SOL: SOL 1

Name of person filing report: HuntJ

Non-nominal systems: All Nominal

Notes on non-nominal systems: N/A

Generator (hours run): N/A (Arrived during day when already off)

Solar— SOC % (Before generator is run at night) 80%

Diesel – 50% Full.

Propane – 80 Percent Volume .

Ethanol Free Gasoline (5 Gallon containers for ATV) – gallons. 5 gallons

Water (trailer) – 1000 Gallons!!!.

Water (static) – 500 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – No

Water Meter: 46 gallons

Toilet tank emptied: No

ATV’s Used: Honda, 300, 350.1, 350.2, 350.3

Oil Added? No

ATV Fuel Used: Negligible (Test drive only)

# Hours the ATVs were Used today: 00:20 hours

Notes on ATVs: They’re running great!!

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? No

General notes and comments: Happy to see the Red Planet!

Summary of internet: So far so good!

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Everything is Awesome!!

Operations Report – December 1st

Crew 183 + Operations Report 1 Dec 2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Phone is still not functional, Low water pressure in Hab

Notes on non-nominal systems: HudgesNet was here all day but unable
to repair the phone. They will be working on it again on Monday. The
filter than normally clogs up the water line in the Hab is missing,
but two other filters were cleaned (Bathroom and kitchen sink) and
water pressure is once again nominal throughout the Hab. A new master
filter will be installed to forever end this problem, which was
probably why the old water heater died.

Generator (hours run): off 8 am on 6 pm

Solar— SOC 94% (Before generator is run at night)

Diesel – 50 %

Propane – 100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons. Curtis took the water trailer to town to fill it.

Water (static) – not known gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not known

Toilet tank emptied: no

ATV’s Used: yes

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 2

Notes on ATVs: nothing to report

Deimos rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR, Operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Not operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? No

General notes and comments: Remember that for the first month of use,
the new rovers will not hold a charge as well as Deimos. We took the
three new ones on a 5 mile round trip and they were each reduced to
about 75%. It takes a while for the batteries to perform at their
peak.

I’d like to say that the rovers were made operational by Brandon and I
but it was actually Chris who did a majority of the work and taught me
a lot about the right tools for the job. Who is Chris? Well, he’s
the guy who was here to fix my phone. It took so long he offered to
help.

The static tank was insulated today. A new heater was installed and
the pipe to the Hab was insulated as well. Shannon’s water tank was
delivered and will be installed tomorrow.

Water heater: After determining that the new water heater would not
fit in the spot where the old one was mounted, it was moved to a spot
on the lower deck just to the right of the ductless propane heater.
It was operational, but would not function correctly until the water
pressure was restored. It is currently connected to the heater’s
propane line. I’m assuming Curtis will reconnect the heater as soon
as he has the supplies.

Summary of internet: not known

Summary of suits and radios: nothing to report

Summary of Hab operations: Water system insulated on the Hab’s
exterior. New hot water heater installed in new location. Water
pressure restored. Water trailer to be refilled.

Summary of GreenHab operations: The thinned plants look mostly good.
One of the tomatoes I tried to transplant does not look good. The
rest look fine. Door was open for a few hours today but it was cold
enough that it stayed mostly warm and not hot.

Summary of ScienceDome operations: Brandon concluded his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Nothing, thank you.

Operations Report – November 30th

Crew 183 + Operations Report 30NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 6 pm

Solar— SOC 90% (Before generator is run at night)

Diesel –  50 %

Propane – 100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 4

Notes on ATVs: nothing to report

Deimos rover used:  yes

Hours:

Beginning charge: 100

Ending charge: 94

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Burn barrel not complete. Turns out you
can’t just poke holes in it with a pickaxe.  We have a plan for
tomorrow.  Also, contractors will be onsite all day tomorrow.

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: nothing to report

Summary of GreenHab operations: Door was opened this afternoon.
Thinned the babies.  Used 4 gallons of water for watering and prepping
the soil for herbs.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Nothing tonight, thank you.

Operations Report – November 29th

Crew 183 + Operations Report 29NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 5:30 pm

Solar— SOC  91% (Before generator is run at night)

Diesel –  50 %

Propane –  100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 3.5

Notes on ATVs: nothing to report

Deimos rover used:  no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments:  Propane was delivered today.  A leak was
found in the line of the small tank at the director’s house, and it
was fixed.  We began work on the second rover today but then decided
to go to storage and see it we could find better tools.  We found a
socket set that was the right size!  We will get back on it tomorrow.
We also brought a 55 gallon drum from town.  Tomorrow I will pound
holes in it and set it up as a burn barrel.  I will also need to clean
up the old burn barrel area as non-burnable garbage was burned and it
will need to be picked up and taken for disposal.  Lucky me!

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: Three new smoke detectors arrived today.

Summary of GreenHab operations: Door was opened this afternoon.  Two
gallons of water was used in the new planters.  I did not water the
plants today.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Thanks for all of
the help with the rover installation question last night.  No
questions tonight.

Operations Report – November 28th

Crew 183 + Operations Report 28NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 6 pm

Solar— SOC  92% (Before generator is run at night)

Diesel –  50 %

Propane –  not known psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –  not known gallons.

Water (trailer) –  0 gallons.

Water (static) –   300 gallons

Trailer to Static Pump used –  no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 3.5

Notes on ATVs: nothing to report

Deimos rover used:  yes

Hours: not recorded (but crews need to record this information)

Beginning charge: 100

Ending charge: 95

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: We began putting the new rovers together
today.  I decided that we would construct the cage, but not modify the
seat at this time.  Instead, we will just remove the back seat rest,
which was what the Mars 160 crew wanted to do originally but I
insisted on the modification.  This will still be done on these but at
a later date because it is much more secure.  Maybe when Yusuke is
here with his crew, since he modified them last time and will be able
to help me to do it.  In the meantime, crews will still be able to use
them.  They will require about a month of short trips before they are
able to go on long ones like Deimos.

Summary of internet:  not known

Summary of suits and radios:  Ashton has recommended that all the old
radios that are sketchy be taken out of service.  I agree.  Four new
radios and headsets will be added to the Hab before the next crew
arrives.  Old sketchy ones will be removed.

Summary of Hab operations: Two new CO alarms were delivered to the
Hab.  We will mount one each on the lower and upper deck before the
new crew arrives.

Summary of GreenHab operations:  Two new planters were added to the
GreenHab.  They are for an herb garden and a greens garden.  I am
prepping the soil.  I turned the heater down to 55 degrees F.  Today I
left the door open all day.  Used 6 gallons of water for plants and
planters.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support:  We have bolts
that have 10.8 and 8.8 on them.  What does that mean in turns of
socket size, if anything?  We don’t have a socket set that fits the
bolts on the rovers, so are trying to figure out if they are metric or
English system.  Any information appreciated.

Operations Report – November 27th

EVA #5 Report

Today our EVA was conducted out-of-sim, with the purpose of further testing our 3D scanning and imagery technology out on the field. Jordan, Ashton, Toruwa, Dean, and I took Deimos and three ATVs to Lith Canyon. We left the hab at approximately 12:20 PM. On the way, we stopped by the location where we obtained the specimens and return them as we found them before.

After spending some time there, we arrived to the Lith Canyon turnout by around 1:15 PM. We walked to the ridge overlooking the canyon and conducted our activities in this area. We stayed here until about 2:30 PM, and made our way back to the Hab by 3:10 PM.

Have a good night Mission Support!

Brandon Ferguson

EVA Officer

Operations Report – November 26th

Crew 183 Operations Report 26-11-2017

SOL: 4

Name of person filing report: Ashton Stoop

Non-nominal systems:

N/A

Notes on non-nominal systems:

Generator (hours run): 14

Solar— 75%

Diesel – 54%

Propane – 40% volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 8 gallons remaining

Water (trailer) – 0 gallons added to static

Water (static) – 300 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – yes

Water Meter: 44

Toilet tank emptied: Yes

ATV’s Used: N/A

Oil Added? No

ATV Fuel Used: 0

# Hours the ATVs were Used today: 0

Notes on ATVs: N/A

Deimos rover used: Yes

Hours: 1

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? NO

General notes and comments:

Summary of internet:

Every evening after dinner time it slows down a lot.

Summary of suits and radios:

A lot of the radios have been giving us issues. Sometimes they work and sometimes they don’t. It seems to be a bit random which is a problem when we are trusting them out on 3-hour EVAs. A few of the suits aren’t charging properly. One of the suits failed today during EVA and we had to rush back to the hab before a crew member’s suit was compromised.

Summary of Hab operations:

Things are operating smoothly.

Summary of GreenHab operations:

Everything is growing well.

Summary of ScienceDome operations:

Sedementation is progressing well.

Summary of RAMM operations:

Not used. It is not yet fully operational.

Summary of health and safety issues:

Waiting and Carbon monoxide detectors.

Questions, concerns and requests to Mission Support:

-There definitely needs to be some further diagnostics performed on suits and radios.

Operations Report – November 25th

Crew 183 Operations Report 25-11-2017

SOL: 3

Name of person filing report: Ashton Stoop

Non-nominal systems:

N/A

Notes on non-nominal systems:

Generator (hours run): 12

Solar— 85%

Diesel – 60%

Propane – 41% volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 8 gallons remaining

Water (trailer) – 0 gallons added to static

Water (static) – 325 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – yes

Water Meter: 48

Toilet tank emptied: No

ATV’s Used: Honda, 350.1

Oil Added? No

ATV Fuel Used: 2 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: Everything ran smoothly today.

Deimos rover used: Yes

Hours: 1

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? NO

General notes and comments:

It’s hot in those EVA suits…

Summary of internet:

Every evening after dinner time it slows down a lot.

Summary of suits and radios:

There have been some issues with radios and headsets over the past few days. We will need to do a test and see which ones are working properly and which aren’t. Some of the EVA suits are having issues. Suit 4 only had one fan working and suit 1 wasn’t working at all. We will double check all the charges to see if that was an issue. One of the XLR cables isn’t working properly. We’re going to measure the current going through and give a more detailed diagnosis in the near future.

Summary of Hab operations:

Things are operating smoothly.

Summary of GreenHab operations:

Everything is growing well and grow lights are operational.

Summary of ScienceDome operations:

Sedementation tests are proceeding as planned. We received a new soil solidity device today.

Summary of RAMM operations:

Not used. It is not yet fully operational.

Summary of health and safety issues:

Waiting and Carbon monoxide detectors.

Questions, concerns and requests to Mission Support:

-Things are going well overall.

Copyright © The Mars Society. All rights reserved. | Main Site